OpenStudy (anonymous):

Ok… Tangent line equations are killing me. How do I find the line tangent to y=(x+3)^1/3 at x=-3 ? Thanks

OpenStudy (anonymous):

not as hard as it looks find the derivative, then evaluate it at \(x=-3\) to find your slope

OpenStudy (anonymous):

since you already have it written in exponential form, the derivative is easy to find \[y=(x+3)^{\frac{1}{3}}\] \[y'=\frac{1}{3}(x+3)^{-\frac{2}{3}}\] but to evaluate this you should put it in radical form as \[y'=\frac{1}{3\sqrt[3]{(x+3)^2}}\]